3(10pt). Solve the initial value problem: y(3) – 5y + 6y = 0, y(0) = 1, y(0) = 0, y (0) = 1.
Answer #1

yol 571+67/20 The arscialloy egri is m?smtómeo mim 5m+6)=0 m(ma2) (n-3) =0 m=0,2,3 The solution is yol) = George og et 100 =70) = 2608438 yol-o =] 2cétigeso 262+36 = 0 alon) = 49,49gedy ytravel > 4ąé+9ge=1 -) 4G2+96-1-)5 by solving 0,@, q= 7/6, CX4